OpenStudy (anonymous):

I need help please: I have been to a "Live Math Tutor" and I am unable to understand the response given. A certain mountain has an elevation of 19,649 ft. In 1915, the glacier on this peak covered 4 acres. By 2000 this glacier had melted to only 1 acre. a) Assume that the glacier melted at a constant rate each year. Find the yearly rate. b) Use answer from part A to write a linear equation that gives acreage A of this glacier t years past 1915. I need to know the following: * the yearly rate of change is ________acres/year? *what is the acreage of the glacier t years after 1915? A=___

OpenStudy (anonymous):

Sorry...made a mistake that last part of the problem should read: What is the equation that gives the acreage of the glacier t years after 1915?

hero (hero):

Well, if we consider only the glacier acreage and don't pay attention to the elevation of the mountain, then we can create the linear equation.

OpenStudy (anonymous):

ok, I am listening...

hero (hero):

Why do we not pay any attention to that? Because we're not asked anything concerning it.

OpenStudy (anonymous):

is all confusing to me

hero (hero):

I know. But, next time, to help you out, you might want to consider creating a relationship chart such as this:

hero (hero):

Often times, for such problems, there will be a relationship between two things. In this case it is Glacier Height vs. Years.

hero (hero):

Which can be plotted on an xy plane or chart.

hero (hero):

If you have two points on the same line, you can figure out the slope and the equation of the line.

OpenStudy (anonymous):

should the yearly rate of change be answered in a fraction such as: 3/86 acres/year?

hero (hero):

Do you know how to find the equation of a line given two points on the line?

OpenStudy (anonymous):

no I do not

OpenStudy (anonymous):

I have been working on this same problem off & on since yesterday...I am trying

hero (hero):

Well, if given two points, you can use the slope formula to find the slope or rate of increase /decrease of a line or relationship between two variables.

hero (hero):

\[m = \frac{y - y_1}{x - x_1}\]

hero (hero):

\[m =\frac{4 - 1}{1915 - 2000} = -\frac{3}{85}\]

OpenStudy (anonymous):

looking, trying 2 learn...ty for helping me

hero (hero):

That's the slope of the line or the rate of decrease of the relationship between the height of the glacier and the amount of years elapsed.

OpenStudy (anonymous):

Well, I guess I was close to the correct answer, but as we know close is incorrect

hero (hero):

Maybe I should have written it like this: \[m = \frac{\Delta\text{Glacier Height}}{\Delta \text{Years}} \]

OpenStudy (anonymous):

So, the yearly rate of change is; -3/85 Acres/year? Am I writing this correctly?

OpenStudy (anonymous):

A negative fraction?

hero (hero):

Yes, that is correct.

hero (hero):

Negative because the glacier height is decreasing over time.

OpenStudy (anonymous):

starting to smile...

hero (hero):

We still need to figure out the linear equation though. The equation of a line is:\[ y = mx + b\]

OpenStudy (anonymous):

Ok, so for the 2nd part...what is the linear equation that gives the acreage of the glacier t years after 1915?

OpenStudy (anonymous):

The way that this is worded is so hard to understand

OpenStudy (anonymous):

y=mx+b...is mx = the glacier height?

hero (hero):

Not to worry, it just means that \(t\) is the variable we use for years rather than \(x\).

hero (hero):

\(y\) is the glacier height \(m\) is the slope of the line or rate of change between y and x \(x\) is the amount of years that goes by

hero (hero):

But in this case, \(x = t\)

hero (hero):

Let's input the slope and one of our points to find b: \[(2000) = -(3/85)(1) + b\] \[2000 = -3/85 + b\]

hero (hero):

Now we just need to solve for b

OpenStudy (anonymous):

y is the glacier height? t is the years?

OpenStudy (anonymous):

oh no, have u left me??

hero (hero):

I haven't left you.

hero (hero):

Solve for b, then we can continue.

OpenStudy (anonymous):

TY so much..LOL

OpenStudy (anonymous):

is b = 86

OpenStudy (anonymous):

(3/85)(1)+b = 86???

hero (hero):

What's 2000 + 3/85 ?

OpenStudy (anonymous):

I am unsure....I feel so dumb..

OpenStudy (anonymous):

2385???

hero (hero):

No

OpenStudy (anonymous):

2000 with a fraction of 3/85??

hero (hero):

Yes

hero (hero):

Anytime you see something like that, you can re-write it as a mixed fraction: \[2000\frac{3}{85}\]

OpenStudy (anonymous):

b = 2000 3/85??

hero (hero):

But we usually don't write b in that form

OpenStudy (anonymous):

ok, I am listening

OpenStudy (anonymous):

When we are done will also explain the hero/medal/fan process 2 me

hero (hero):

Well, basically, what you do is sort of round off. 3/85 is closer to zero. So \(b \approx 2000\)

OpenStudy (anonymous):

so b is rounded off to the closest 0. b = 2000:)

OpenStudy (anonymous):

A=3t/85 + 2000? This is the linear equation that gives the acreage of the glacier t years after 1915?

OpenStudy (anonymous):

u there??

hero (hero):

But wait...that slope should be negative.

hero (hero):

\[A = 2000 - \frac{3}{85}t\]

OpenStudy (anonymous):

So, that is my answer...A=2000-3/85 t ??? this is the linear equation that I am looking for?

hero (hero):

Yes

OpenStudy (anonymous):

The yearly rate of change is -3/85 & we are using t instead of x in the given formula

OpenStudy (anonymous):

Ok, 1 last thing how do I commend u...MY HERO

OpenStudy (anonymous):

I am new to this new friend!!

OpenStudy (anonymous):

can i log in again & ask additional questions?

hero (hero):

You can ask as many questions as you like.

OpenStudy (anonymous):

how do I commend u & recognize u for your WONDERFUL EFFORTS

hero (hero):

That's not really necessary.

OpenStudy (anonymous):

ok, I REALLY APPRECIATE YOUR TIME & EFFORTS....MAYBE I could mail u a ham or something nice??? lol

hero (hero):

Mail me a ham? That would be rather odd. I don't usually receive hams in the mail

hero (hero):

Let me check something real quick.

OpenStudy (anonymous):

LOL I have never mailed a ham.....TY

hero (hero):

Okay, yep, the result has been confirmed correct.